OpenStudy (anonymous):

(x+3) (x+3)^3 simplify

OpenStudy (anonymous):

(a+b)^3 = a^3 + b^3 + 3ab(a+b)

OpenStudy (anonymous):

see there are two ways...... either expand the cube and multiply/.... or use binomial expansion (x+3)^4.

mathslover (mathslover):

\[\large{(x+3)(x+3)^3}\] \[\large{(a+b)^3=a^3+b^3+3ab(a+b)->identity}\]

mathslover (mathslover):

@cdelomas have u learnt binomial theorem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

binomial theorem or pascals triangle will work

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

http://www.mathsisfun.com/pascals-triangle.html

OpenStudy (anonymous):

i will check it out and comeback

mathslover (mathslover):

forget binomial theorem : \[\large{(x+3)(x+3)^3=(x+3){(x+3)(x+3)(x+3)}}\] \[\large{(x+3)^2(x+3)^2}\] \[\large{(x+3)^2=x^2+3^2+2(x)(3)}\] \[\large{x^2+9+6x}\] \[\large{(x^2+9+6x)(x^2+9+6x)}\] \[\large{x^4+9x^2+6x^3+9x^2+81+54x+6x^3+54x+36x^2}\] \[\large{x^4+12x^3+54x^2+108x+81}\]

mathslover (mathslover):

got it ?